# Unsupervised Meta-Learning via In-Context Learning

### v2

@@ A.8 Quantitative analysis of learning phases
Figures 7(a) and 7(b) show the results for CAMeLU on the miniImageNet and CUB datasets, respectively. For miniImageNet, we obtain the approximation function to be $f(x)=0.04+\frac{0.54}{1+9636e^{-0.43x}}$. Moreover, the number of epochs where the learning phase begins is $15$ and the number of epochs where the generalization phase begins is $29$. On the other hand, for CUB dataset, the approximation function is $f(x)=0.01+\frac{0.48}{1+25530e^{-0.47x}}$, and the number of epochs where the following phases begin is $16$ and $28$.
Figures 7(c) and 7(d) show the results for CAML on the miniImageNet and the CUB datasets, respectively. Similarly, we obtain the approximation function for the miniImageNet to be $f(x)=0.06+\frac{0.56}{1+1326e^{-0.25x}}$. and the number of epochs where the learning phase and the generalization phase begin is $18$ and $42$. Finally, for the CUB dataset, the approximation function is $f(x)=-0.07+\frac{0.50}{1+648e^{-0.13x}}$, and the number of epochs where the following phases begin is $29$ and $74$.
@@ A.10 Complete results with standard deviations
Table 10: Performance comparison on miniImageNet, CIFAR-fs, CUB, Aircraft, and Meta-iNat datasets for 5-way 1-shot and 5-way 5-shot scenarios. Cross-domain approaches are trained using ImageNet-964 and a ResNet-50 feature extractor. The symbol $\dagger$ indicates results that are affected by data leakage. The bold font highlights the best performing UML approach for each setting. Results show the mean and standard deviations across three complete runs of the algorithms. This table refers to Table 1 in Section 4.3.
@@
Table 11: Accuracy results obtained training PsCo, BECLR, and CAMeLU with a small-scale dataset, namely miniImageNet, denoted as (mini) in the table. Results show both in-domain performance (on the test set of miniImageNet) and cross-domain performance on CIFAR-fs, CUB, Aircraft, and Meta-iNat. The mean and standard deviation across three complete runs of the algorithms. This table refers to Table 2 in Section 4.5.
@@
Table 12: Comparison between CAMeLU and SSL approaches for the 5-way 1-shot and 5-way 5-shot scenario on miniImageNet, CIFAR-fs, CUB, Aircraft, and Meta-iNat. The symbol $\dagger$ indicates results that are affected by data leakage. Results show the mean and standard deviations across three complete runs of the algorithms. This table refers to Table 3 in Section 4.6.

### v3

@@ A.8 Quantitative analysis of learning phases
Figures 8(a) and 8(b) show the results for CAMeLU on the miniImageNet and CUB datasets, respectively. For miniImageNet, we obtain the approximation function to be $f(x)=0.04+\frac{0.54}{1+9636e^{-0.43x}}$. Moreover, the number of epochs where the learning phase begins is $15$ and the number of epochs where the generalization phase begins is $29$. On the other hand, for CUB dataset, the approximation function is $f(x)=0.01+\frac{0.48}{1+25530e^{-0.47x}}$, and the number of epochs where the following phases begin is $16$ and $28$.
Figures 8(c) and 8(d) show the results for CAML on the miniImageNet and the CUB datasets, respectively. Similarly, we obtain the approximation function for the miniImageNet to be $f(x)=0.06+\frac{0.56}{1+1326e^{-0.25x}}$. and the number of epochs where the learning phase and the generalization phase begin is $18$ and $42$. Finally, for the CUB dataset, the approximation function is $f(x)=-0.07+\frac{0.50}{1+648e^{-0.13x}}$, and the number of epochs where the following phases begin is $29$ and $74$.
@@ A.10 Complete results with standard deviations
Table 12: Performance comparison on miniImageNet, CIFAR-fs, CUB, Aircraft, and Meta-iNat datasets for 5-way 1-shot and 5-way 5-shot scenarios. Cross-domain approaches are trained using ImageNet-964 and a ResNet-50 feature extractor. The symbol $\dagger$ indicates results that are affected by data leakage. The bold font highlights the best performing UML approach for each setting. Results show the mean and standard deviations across three complete runs of the algorithms. This table refers to Tab. 1 in Sect. 4.3.
@@
Table 13: Accuracy results obtained training PsCo, BECLR, and CAMeLU with a small-scale dataset, namely miniImageNet, denoted as (mini) in the table. Results show both in-domain performance (on the test set of miniImageNet) and cross-domain performance on CIFAR-fs, CUB, Aircraft, and Meta-iNat. The mean and standard deviation across three complete runs of the algorithms. This table refers to Tab. 2 in Sect. 4.5.
@@
Table 14: Comparison between CAMeLU and SSL approaches for the 5-way 1-shot and 5-way 5-shot scenario on miniImageNet, CIFAR-fs, CUB, Aircraft, and Meta-iNat. The symbol $\dagger$ indicates results that are affected by data leakage. Results show the mean and standard deviations across three complete runs of the algorithms. This table refers to Tab. 3 in Sect. 4.6.